OpenStudy (abbycross167):

Can someone please help me answer this math question? Please? I'll fan and give a medal!! Question - Factor using a method that works best showing the steps - 125a^4 – 5a^2b2 (Hint: Try GCF first, then see if the result can be factored further.)

OpenStudy (abbycross167):

@jim_thompson5910 could you try to help me with this question? please?

OpenStudy (bmk614):

What do you get as the GCF?

OpenStudy (abbycross167):

1?

OpenStudy (bmk614):

To take the GCF you need to find the biggest common numbers or most amount of variables. You would start with constants, then move onto variables. I get 5a^2.

OpenStudy (bmk614):

Do you understand how to do that?

OpenStudy (abbycross167):

to be honest I don't really understand ... the b2 isn't b^2

OpenStudy (abbycross167):

Sir/ma'am, I don't understand how you got 5a^2

OpenStudy (bmk614):

Well out of the 125 and 10 (same as 5*2), since they are like terms, the GCF is 5. The GCF of a^4 and a^2 would be a^2. Since there is no b in the first number, the GCF does not exist.

OpenStudy (bmk614):

You multiply the 5 and a^2 together, which gives you your GCF.

OpenStudy (abbycross167):

So step 1 would be Step 1: 5*a^2

OpenStudy (bmk614):

Correct. Then you would take that out of each section. You are, in a sense, backwards distributing.

OpenStudy (abbycross167):

okay umm, would you mind showing me this step? I don't understand what you mean but I thinnk i'll understand if you show me

OpenStudy (bmk614):

Well you divide both parts, by the GCF. You cross off like terms.

OpenStudy (abbycross167):

so like this ? 125a^4 – 5a^2b2 125/10 = 12.5 5/10 = 0.5 2/10=5 ?

OpenStudy (bmk614):

No. Hold on.

OpenStudy (bmk614):